OpenStudy (anonymous):

solve using the multiplication principle 9x = -72

OpenStudy (aroub):

x=-8

OpenStudy (anonymous):

9 x = -8 * 9 x = -8

OpenStudy (aroub):

9x=-72*1/9 x=-8